### v2

# Flow-GRPO: Training Flow Matching Models via Online RL

Abstract
We propose Flow-GRPO, the first method integrating online reinforcement learning (RL) into flow matching models. Our approach uses two key strategies: (1) an ODE-to-SDE conversion that transforms a deterministic Ordinary Differential Equation (ODE) into an equivalent Stochastic Differential Equation (SDE) that matches the original model’s marginal distribution at all timesteps, enabling statistical sampling for RL exploration; and (2) a Denoising Reduction strategy that reduces training denoising steps while retaining the original inference timestep number, significantly improving sampling efficiency without performance degradation. Empirically, Flow-GRPO is effective across multiple text-to-image tasks. For complex compositions, RL-tuned SD3.5 generates nearly perfect object counts, spatial relations, and fine-grained attributes, boosting GenEval accuracy from $63\%$ to $95\%$. In visual text rendering, its accuracy improves from $59\%$ to $92\%$, significantly enhancing text generation. Flow-GRPO also achieves substantial gains in human preference alignment. Notably, very little reward hacking occurred, meaning rewards did not increase at the cost of appreciable image quality or diversity degradation.
Figure 1: (a) GenEval performance rises steadily throughout Flow-GRPO’s training and outperforms GPT-4o. (b) Image quality metrics on DrawBench [1] remain essentially unchanged. (c) Human Preference Scores on DrawBench improves after training. Results show that Flow-GRPO enhances the desired capability while preserving image quality and exhibiting minimal reward-hacking.
1 Introduction
Flow matching [2, 3] models have become dominant in image generation [4, 5] due to their solid theoretical foundations and strong performance in producing high-quality images. However, they often struggle with composing complex scenes involving multiple objects, attributes, and relationships [6, 7], as well as text rendering [8]. At the same time, online reinforcement learning (RL) [9] has proven highly effective in enhancing the reasoning capabilities of large language models (LLMs) [10, 11]. While previous research has mainly focused on applying RL to early diffusion-based generative models [12] and offline RL techniques like direct preference optimization [13] for flow-based generative models [14, 15], the potential of online RL in advancing flow matching generative models remains largely unexplored. In this study, we explore how online RL can be leveraged to effectively improve flow matching models.
Training flow models with RL presents several critical challenges: (1) Flow models rely on a deterministic generative process based on ordinary differential equations (ODEs) [3], meaning they cannot sample stochastically during inference. In contrast, RL relies on stochastic sampling to explore environments, where the agent learns by trying different actions and receiving feedback. This need for stochasticity in RL conflicts with the deterministic nature of flow matching models. (2) Online RL requires efficient sampling to gather training data, but flow models typically involve many iterative steps to generate each sample, significantly reducing sampling efficiency. This issue becomes even more critical with large, advanced models [5, 4], which are computationally intensive. Therefore, to make RL effective in tasks such as image or video generation, improving sampling efficiency within RL training is crucial.
To address these challenges, we propose Flow-GRPO, which integrates GRPO [16] into flow matching models for text-to-image (T2I) generation, using two key strategies. First, we adopt the ODE-to-SDE strategy to overcome the deterministic nature of the original flow model. By converting the ODE-based flow into an equivalent Stochastic Differential Equation (SDE) framework, we introduce randomness while preserving the original marginal distributions. Second, to improve sampling efficiency in online RL, we apply the Denoising Reduction strategy—reducing denoising steps during training while keeping the full schedule during inference. Our experiments show that using fewer steps maintains performance while significantly reducing data generation costs.
We evaluate Flow-GRPO on T2I tasks with various reward types. (1) Verifiable rewards, using the GenEval [17] benchmark and visual text rendering task. GenEval includes compositional image generation tasks (e.g., generating specific object counts, colors, and spatial relationships), which can be automatically assessed with traditional detection methods. Flow-GRPO improves the accuracy of Stable Diffusion 3.5 Medium (SD3.5-M) [4] from 63% to 95% on GenEval, outperforming the state-of-the-art GPT-4o [18] model. For visual text rendering, SD3-M’s accuracy increases from 59% to 92%, greatly enhancing its text generation ability. (2) Model-based rewards, such as the human preference Pickscore [19] reward, show that our framework is task independent, demonstrating its generalizability and robustness. Importantly, all improvements are achieved with very little reward hacking, as demonstrated in Figure 1.
To summarize, the contributions of Flow-GRPO are as follows:
  • We are the first to introduce GRPO to flow matching models by converting deterministic ODE sampling into SDE sampling, showing that online RL is highly effective for T2I tasks. Specifically, our Flow-GRPO improves SD3.5-Medium accuracy from 63% to 95% without compromising image quality.
  • We find that online RL for flow matching models can use fewer denoising steps for training sample generation, significantly accelerating training. Full denoising steps are still used during testing to maintain performance.
  • We show that the Kullback-Leibler (KL) constraint is effective in preventing reward hacking, where reward increases but image quality or diversity decreases. We also emphasize that KL regularization is not empirically equivalent to early stopping. By adding an appropriate KL, we can achieve the same high reward as the KL-free version while maintaining image quality, though with a longer training time.
2 Related Work
RL for LLM.
Online RL has proven effective in boosting the reasoning abilities of large language models (LLMs), notably DeepSeek-R1 [10] and OpenAI-o1 [11]. By iteratively alternating between exploration and exploitation, these systems generate increasingly detailed responses and steadily enhance overall performance. Most successful pipelines employ policy gradient algorithms such as Proximal Policy Optimization (PPO) [20] and the value-free Group Relative Policy Optimization (GRPO) [16]. Because GRPO dispenses with a separate value network, it is more memory efficient; hence we adopt GRPO in this work. Nonetheless, PPO could be incorporated into a flow matching model through an analogous procedure.
Diffusion and Flow Matching.
Diffusion-based [21, 22, 23] generative models create a forward Markov chain that incrementally adds Gaussian noise to data and train a neural network to approximate the reverse-time denoising vector field; integrating this learned field by discrete DDPM steps or probability flow SDE solvers can yields high fidelity samples. Flow matching [2, 3] instead treats generation as learning a continuous-time normalizing flow whose vector field is trained by directly matching the velocity, enables deterministic sampling in a few ODE steps. Thanks to its efficiency, flow matching attains competitive FID with orders-of-magnitude fewer denoising steps than diffusion samplers and has already become the dominant approach in current image generation [4, 5] and video generation [24, 25, 26, 27] models.  [28, 29] unifies flow-based and diffusion-based methods using the SDE and ODE perspective. Our work builds on their theoretical foundations and introduce GRPO to flow-based models.
Alignment for T2I.
Recent work on aligning pretrained T2I models with human preferences has pursued several complementary directions. The main approaches are: (1) direct fine-tuning with scalar reward signals [30, 31, 32, 33]; (2) Reward-Weighted Regression (RWR) [34, 35, 36]; (3) Direct Preference Optimization (DPO) and related variants,[37, 38, 39, 14, 40, 41, 42, 43, 44, 36]; (4) Proximal Policy Optimization (PPO)-style policy gradients [45, 46, 47, 48, 49]; and (5) training-free alignment techniques [50, 51, 52]. These methods have successfully aligned T2I models with human preferences, improving aesthetics and semantic consistency. Building on this progress, we introduce GRPO for flow-matching models—the backbone of today’s state-of-the-art T2I systems. Concurrent work [53] apply GRPO to text-to-speech flow models, but instead of converting the ODE to an SDE to inject stochasticity, they reformulate velocity prediction as estimating a Gaussian distribution (predicting both the mean and variance of velocity), which necessitates retraining the pretrained model. [54] also recognize that stochasticity can be introduced from an SDE perspective, yet their work concentrates on inference-time scaling.
3 Preliminaries
In this section, we introduce the mathematical formulation of flow matching (Section 3.1) and describe how the denoising process can be mapped as a multi-step MDP (Section 3.2).
3.1 Flow Matching
Let ${\bm{x}}_{0}\sim X_{0}$ denote a data sample drawn from the true data distribution, and ${\bm{x}}_{1}\sim X_{1}$ denote a noise sample. Recent advanced image-generation models (e.g., [4, 5]) and video-generation models (e.g., [24, 26, 25, 27]) adopt the Rectified Flow [3] framework, which defines the “noised” data ${\bm{x}}_{t}$ as
${\bm{x}}_{t}=(1-t)\,{\bm{x}}_{0}\;+\;t\,{\bm{x}}_{1},$ (1)
for $t\in[0,1]$. Then a transformer model are trained to directly regress the velocity field ${\bm{v}}_{\theta}({\bm{x}}_{t},t)$ by minimizing the Flow Matching objective [2, 3]:
$\mathcal{L}(\theta)=\mathbb{E}_{t,\;{\bm{x}}_{0}\sim X_{0},\;{\bm{x}}_{1}\sim X_{1}}\bigl{[}\;\|{\bm{v}}\;-\;{\bm{v}}_{\theta}({\bm{x}}_{t},t)\|^{2}\bigr{]},$ (2)
where the target velocity field is ${\bm{v}}={\bm{x}}_{1}-{\bm{x}}_{0}$.
Figure 2: Overview of Flow-GRPO. Given a prompt set, we introduce an ODE-to-SDE strategy to enable stochastic sampling for online RL. With Denoising Reduction (only T = 10 steps), we efficiently gather low-quality but still informative trajectories. Rewards from these trajectories feed the GRPO loss, which updates the model online and yields an aligned policy.
3.2 Denoising as an MDP
According to [12], the iterative denoising process in diffusion models can be framed as an Markov decision process (MDP) $(\mathcal{S},\mathcal{A},\rho_{0},P,R)$. The state at time step p $t$ is ${\bm{s}}_{t}\!\triangleq\!({\bm{c}},t,{\bm{x}}_{t})$, the action is the denoised sample ${\bm{a}}_{t}\!\triangleq\!{\bm{x}}_{t-1}$ produced by the network, and the policy is the conditional model $\pi({\bm{a}}_{t}\mid{\bm{s}}_{t})\!\triangleq\!p_{\theta}({\bm{x}}_{t-1}\mid{\bm{x}}_{t},{\bm{c}})$. The transition kernel is deterministic $P({\bm{s}}_{t+1}\mid{\bm{s}}_{t},{\bm{a}}_{t})\!\triangleq\!(\delta_{\bm{c}},\delta_{t-1},\delta_{{\bm{x}}_{t-1}})$ and the initial-state distribution is $\rho_{0}({\bm{s}}_{0})\!\triangleq\!(p({\bm{c}}),\delta_{T},\mathcal{N}(\mathbf{0},\mathbf{I}))$, in which $\delta_{y}$ is the Dirac delta distribution with nonzero density only at $y$. The reward is non-zero only at the final step, $R({\bm{s}}_{t},{\bm{a}}_{t})\!\triangleq\!r({\bm{x}}_{0},{\bm{c}})$ if $t\!=\!0$ and 00 otherwise.
4 Flow-GRPO
In this section, we present the Flow-GRPO method, a GRPO-inspired algorithm that improves flow models through online RL. We begin by revisiting the core concept of GRPO [16] and recasting it in the context of flow matching. Next, we discuss how to convert the deterministic probability-flow ODE sampler into a reverse-time SDE with equivalent marginal distribution, thereby introducing stochasticity into the generation process—an essential ingredient for applying GRPO or other RL techniques. Finally, we introduce Denoise Reduction (Section 4.3), a practical data sampling method that dramatically accelerates training without degrading performance.
4.1 GRPO on Flow Matching
Reinforcement learning aims to learn a policy that maximizes expected cumulative reward under a given environment. Typically the goal is to optimize a policy $\pi_{\theta}$ with a regularized objective of the form:
$\max_{\theta}\mathbb{E}_{({\bm{s}}_{0},{\bm{a}}_{0},\dots,{\bm{s}}_{T},{\bm{a}}_{T})\sim\pi_{\theta}}\Bigg{[}\sum_{t=0}^{T}\Bigg{(}R({\bm{s}}_{t},{\bm{a}}_{t})-\beta D_{\text{KL}}(\pi_{\theta}(\cdot\mid{\bm{s}}_{t})||\pi_{\text{ref}}(\cdot\mid{\bm{s}}_{t}))\Bigg{)}\Bigg{]}.$ (3)
Among various RL algorithms, GRPO [16] provides a lightweight alternative to policy based methods such as PPO [20]. Unlike PPO [20], which requires training a separate value function—normally a separate model of similar size to the policy network, GRPO introduces a group relative formulation to estimate the advantage. In our setting, we recast the GRPO framework into the context of flow matching. Recall that the denoising process can be formulated as an MDP, as shown in Section 3.2. Given a prompt ${\bm{c}}$, the flow model $p_{\theta}$ samples a group of $G$ individual images $\{{\bm{x}}^{i}_{0}\}_{i=1}^{G}$ and the corresponding reverse-time trajectories $\{({\bm{x}}^{i}_{T},{\bm{x}}^{i}_{T-1},\cdots,{\bm{x}}^{i}_{0})\}_{i=1}^{G}$. Then, the advantage of the $i$-th image is calculated by normalizing the group-level rewards as follows:
$\hat{A}^{i}_{t}=\frac{R({\bm{x}}^{i}_{0},{\bm{c}})-\text{mean}(\{R({\bm{x}}^{i}_{0},{\bm{c}})\}_{i=1}^{G})}{\text{std}(\{R({\bm{x}}^{i}_{0},{\bm{c}})\}_{i=1}^{G})}.$ (4)
GRPO employs a clipped objective of Eq. 3, together with a KL penalty term:
$\mathcal{J}_{\text{Flow-GRPO}}(\theta)$ $=\mathbb{E}_{{\bm{c}}\sim\mathcal{C},\{{\bm{x}}^{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid{\bm{c}})}$ (5)
$\Bigg{[}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{T}\sum_{t=0}^{T-1}\Bigg{(}\min\Big{(}r^{i}_{t}(\theta)\hat{A}^{i}_{t},\ \text{clip}\Big{(}r^{i}_{t}(\theta),1-\varepsilon,1+\varepsilon\Big{)}\hat{A}^{i}_{t}\Big{)}-\beta D_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\Bigg{)}\Bigg{]},$
where
$r^{i}_{t}(\theta)=\frac{p_{\theta}({\bm{x}}^{i}_{t-1}\mid{\bm{x}}^{i}_{t},{\bm{c}})}{p_{\theta_{\text{old}}}({\bm{x}}^{i}_{t-1}\mid{\bm{x}}^{i}_{t},{\bm{c}})}.$ (6)
4.2 ODE-to-SDE
In Eq. 4 and Eq. 6, GRPO relies on a stochastic sampling procedure in order to generate a diverse batch of trajectories for advantage estimation and policy exploration. In the case of diffusion models, the forward process consists of a sequence of random Gaussian noising steps, and its reverse-time sampler is implemented via a Markov chain of Gaussians with progressively decreasing variance—equivalently, an approximate solver for the score-based SDE. Thus, diffusion sampling is naturally stochastic.
While in the case of probability flow model, as previously discussed in Eq. 1, its forward process can be formulated as a deterministic ODE:
$\mathrm{d}{\bm{x}}_{t}={\bm{v}}_{t}\mathrm{d}t,$ (7)
where ${\bm{v}}_{t}$ is modeled via the flow matching objective in Eq. 2. A common sampling method is to discretize this ODE (e.g., via the explicit Euler method or a higher-order integrator) yielding a one-to-one mapping between successive time steps. For example, with a fixed step size $\Delta t$, the Euler scheme gives:
${\bm{x}}_{t-1}={\bm{x}}_{t}+\Delta t{\bm{v}}_{\theta}({\bm{x}}_{t},t,{\bm{c}})$ (8)
which produces a single, deterministic trajectory $\{{\bm{x}}_{T},{\bm{x}}_{T-1},\cdots,{\bm{x}}_{0}\}$ given any initial ${\bm{x}}_{T}$ and prompt ${\bm{c}}$.
This deterministic approach fails to meet GRPO’s policy update requirements in two key ways: (1) Eq. 6 requires computing $p({\bm{x}}_{t-1}\mid{\bm{x}}_{t},{\bm{c}})$, but deterministic sampling cannot provide probabilities. This limitation affects not only GRPO but all policy gradient methods, including the basic online REINFORCE [55] policy gradient, which also require probability computation. (2) Exploration is crucial for RL performance. As shown in Section 5.3, reduced randomness significantly lowers training efficiency. Deterministic sampling, which lacks randomness beyond the initial seed, is especially problematic.
To address this limitation, we convert the deterministic Flow-ODE from Eq. 7 into an equivalent SDE that matches the original model’s marginal probability density function at all timesteps. We outline the key process here. A detailed proof is provided in Appendix A. Following [23, 28, 29], we construct a reverse-SDE formulation that preserves the marginal distribution:
$\mathrm{d}{\bm{x}}_{t}=\left({\bm{v}}_{t}({\bm{x}}_{t})-\frac{\sigma_{t}^{2}}{2}\nabla\log p_{t}({\bm{x}}_{t})\right)\mathrm{d}t+\sigma_{t}\mathrm{d}{\bm{w}},$ (9)
where $\mathrm{d}{\bm{w}}$ denotes Wiener process increments and $\sigma_{t}$ control the level of stachasticity during generation. For rectified flow, the marginal score has a relation with velocity:
$\nabla\log p_{t}({\bm{x}})=-\frac{{\bm{x}}}{t}-\frac{1-t}{t}{\bm{v}}_{t}({\bm{x}}).$ (10)
Substituting Eq. 10 into Eq. 9 gives the final SDE:
$\mathrm{d}{\bm{x}}_{t}=\left[{\bm{v}}_{t}({\bm{x}}_{t})+\frac{\sigma_{t}^{2}}{2t}\left({\bm{x}}_{t}+(1-t){\bm{v}}_{t}({\bm{x}}_{t})\right)\right]\mathrm{d}t+\sigma_{t}\mathrm{d}{\bm{w}}.$ (11)
Applying Euler-Maruyama discretization yields the final update rule:
${\bm{x}}_{t+\Delta t}={\bm{x}}_{t}+\left[{\bm{v}}_{\theta}({\bm{x}}_{t},t)+\frac{\sigma_{t}^{2}}{2t}\big{(}{\bm{x}}_{t}+(1-t){\bm{v}}_{\theta}({\bm{x}}_{t},t)\big{)}\right]\Delta t+\sigma_{t}\sqrt{\Delta t}\,\epsilon$ (12)
where $\epsilon\sim\mathcal{N}(0,{\bm{I}})$ injects stochasticity. We use $\sigma_{t}=a\sqrt{\frac{t}{1-t}}$ in this paper, where $a$ is a scalar hyper-parameter that controls the noise level (See Section 5.3 for its impact on performance).
Eq. 12 reveals that the policy $\pi_{\theta}({\bm{x}}_{t-1}\mid{\bm{x}}_{t},{\bm{c}})$ is an isotropic Gaussian distribution. We can easily compute the KL divergence between $\pi_{\theta}$ and the reference policy $\pi_{\text{ref}}$ as a closed form:
$D_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})=\frac{\|\overline{{\bm{x}}}_{t+\Delta t,\theta}-\overline{{\bm{x}}}_{t+\Delta t,\text{ref}}\|^{2}}{2\sigma_{t}^{2}\Delta t}=\frac{\Delta t}{2}\left(\frac{\sigma_{t}(1-t)}{2t}+\frac{1}{\sigma_{t}}\right)^{2}\|{\bm{v}}_{\theta}({\bm{x}}_{t},t)-{\bm{v}}_{\text{ref}}({\bm{x}}_{t},t)\|^{2}$ (13)
4.3 Denoising Reduction
To produce high-quality images, flow models typically require many denoising steps, making data collection costly for online RL. However, we find that large timesteps are unnecessary during online RL training. We can use significantly fewer denoising steps during sample generation, while retaining the original denoising steps during inference to get high-quality samples. Note that we set the timestep $T$ as $10$ in training, while the inference timestep $T$ is set as the original default setting ($T=40$) for SD3.5-M. Our experiments reveals that this approach enables fast training without sacrificing image quality at test time.
5 Experiments
In this section, we empirically evaluate Flow-GRPO’s ability to improving flow matching models on three representative tasks. (1) Composition Image Generation: This task requires precise object arrangement and attribute control. We report the results on GenEval. (2) Visual Text Rendering: a rule-based task that stresses the accurate rendering of the text specified in the prompt. (3) Human Preference Alignment: This task aims to align T2I models with human preferences. We use PickScore [19] as our reward model.
5.1 Experimental Setup
We introduce three tasks, detailing their respective prompts and reward definitions. For hyperparameter details and compute resource specifications, please refer to Appendix B.2 and Appendix B.2.1.
Compositional Image Generation
[…]
  • Counting: $r=|N_{\text{gen}}-N_{\text{ref}}|/N_{\text{ref}}$.
  • Position / Color: If the object count is correct, a partial reward is assigned; the remainder is granted when the predicted position or color is also correct.
Table 1: GenEval Result. Best scores are in blue, second-best in green. Results for models other than SD3.5-M are from [7] or their original papers. Obj.: Object; Attr.: Attribution.
Model Overall Single Obj. Two Obj. Counting Colors Position Attr. Binding
Diffusion Models
LDM [56] 0.37 0.92 0.29 0.23 0.70 0.02 0.05
SD1.5 [56] 0.43 0.97 0.38 0.35 0.76 0.04 0.06
SD2.1 [56] 0.50 0.98 0.51 0.44 0.85 0.07 0.17
SD-XL [57] 0.55 0.98 0.74 0.39 0.85 0.15 0.23
DALLE-2 [58] 0.52 0.94 0.66 0.49 0.77 0.10 0.19
DALLE-3 [59] 0.67 0.96 0.87 0.47 0.83 0.43 0.45
Autoregressive Models
Show-o [60] 0.53 0.95 0.52 0.49 0.82 0.11 0.28
Emu3-Gen [61] 0.54 0.98 0.71 0.34 0.81 0.17 0.21
JanusFlow [62] 0.63 0.97 0.59 0.45 0.83 0.53 0.42
Janus-Pro-7B [63] 0.80 0.99 0.89 0.59 0.90 0.79 0.66
GPT-4o [18] 0.84 0.99 0.92 0.85 0.92 0.75 0.61
Flow Matching Models
FLUX.1 Dev [5] 0.66 0.98 0.81 0.74 0.79 0.22 0.45
SD3.5-L [4] 0.71 0.98 0.89 0.73 0.83 0.34 0.47
SANA-1.5 4.8B [64] 0.81 0.99 0.93 0.86 0.84 0.59 0.65
SD3.5-M [4] 0.63 0.98 0.78 0.50 0.81 0.24 0.52
SD3.5-M+Flow-GRPO 0.95 1.00 0.99 0.95 0.92 0.99 0.86
Figure 3: Qualitative Comparison on the GenEval Benchmark. Our approach demonstrates superior performance in Counting, Colors, Attribute Binding, and Position.
Table 2: Performance on Compositional Image Generation, Visual Text Rendering, and Human Preference benchmarks, evaluated by task performance, image quality, and preference scores. Acc: Accuracy; ImgRwd: ImageReward; UniRwd: UnifiedReward.
Model Task Metric Image Quality Preference Score
GenEval OCR Acc. PickScore Aesthetic DeQA ImgRwd PickScore UniRwd
SD3.5-M 0.63 0.59 21.72 5.39 4.07 0.87 22.34 3.33
Compositional Image Generation
Flow-GRPO (w/o KL) 0.95 4.93 2.77 0.44 21.16 2.94
Flow-GRPO (w/ KL) 0.95 5.25 4.01 1.03 22.37 3.51
Visual Text Rendering
Flow-GRPO (w/o KL) 0.93 5.13 3.66 0.58 21.79 3.15
Flow-GRPO (w/ KL) 0.92 5.32 4.06 0.95 22.44 3.42
Human Preference Alignment
Flow-GRPO (w/o KL) 23.41 6.15 4.16 1.24 3.57
Flow-GRPO (w/ KL) 23.31 5.92 4.22 1.28 3.66
Visual Text Rendering [8]
Text is common in images such as posters, book covers, and memes, so the ability to place accurate and coherent text inside the generated images is crucial for T2V models. In our settings, we define an OCR task, where each prompt follows the template ‘‘A sign that says ‘‘text’’. Specifically, the placeholder ‘‘text’’ is the exact string that should appear in the image. We use GPT to produce 20K training prompts and 1K test prompts. Following [65], we measure text fidelity with the reward
$r=\frac{|N_{\text{gen}}-N_{\text{ref}}|}{N_{\text{ref}}},$
where $N_{\text{gen}}$ is the number of characters detected in the image and $N_{\text{ref}}$ is the number of characters inside the quotation marks in the prompt. This reward also serves as our metric of text accuracy.
Human Preference Alignment [19]
This task aims to align T2I models with human preferences. We use PickScore  [^2] https://huggingface.co/yuvalkirstain/PickScore_v1 as our reward model, which is based on large-scale human annotated pairwise comparisons of images generated from the same prompt. For each image and prompt pair, PickScore provides an overall score that evaluates multiple criteria, such as the alignment of the image with the prompt and its visual quality.
Image Quality Evaluation Metric
Because the T2I model is trained to maximise a predefined reward signal in RL, it is susceptible to reward hacking—situations in which the numerical reward rises but the quality or diversity of the generated images declines. This study aims to make online RL truly effective for T2I generation without compromising quality or diversity. To detect potential reward hacking beyond task-specific accuracy, we evaluate four automatic image quality metrics:
  • Aesthetic score [66]: a CLIP-based linear regressor that predicts an image’s aesthetic score.
  • DeQA score [67]: a multimodal large language model based image-quality assessment (IQA) model that quantifies how distortions, texture damage, and other low-level artefacts affect perceived quality.
  • ImageReward [32]: a general purpose T2I human preference reward model that captures text–image alignment, visual fidelity, and harmlessness.
  • UnifiedReward [68]: a recently proposed unified reward model for multimodal understanding and generation that currently achieves state-of-the-art performance on the human preference assessment leaderboard.
All metrics are evaluated on DrawBench [1], a comprehensive and challenging benchmark with diverse prompts for T2I models.
5.2 Main Results
Figure 1 and Table 1 show Flow-GRPO’s GenEval performance steadily improving during training, ultimately outperforming GPT-4o. This occurs while maintaining both image quality metrics and preference scores on DrawBench, a benchmark with diverse and comprehensive prompts for evaluating general model capabilities. Figure 3 offers qualitative comparisons. Beyond Compositional Image Generation, Table 2 details evaluations on Visual Text Rendering and Human Preference tasks. Flow-GRPO improved text rendering ability, again without decreasing image quality metrics and preference scores on DrawBench. See Figures 8 and 9 in Appendix C.3 for related qualitative examples. For the Human Preference task, image quality did not decrease without KL regularization. However, we found that omitting KL caused a collapse in visual diversity—a form of reward hacking discussed further in Section 5.3. These results demonstrate that Flow-GRPO boosts desired capabilities while causing very little degradation to image quality or visual diversity.
5.3 Analysis
Reward Hacking
We explored two approaches to mitigate reward hacking: (1) Multi-reward ensemble: We combined preference-based rewards (e.g., UnifiedReward, PickScore, ImageReward), image quality rewards (e.g., Aesthetic Score, DeQA) with task reward. However, this also led to local blurriness and reduced diversity. Due to the challenge of covering all aspects of generation with diverse reward models, we abandoned this approach. (2) KL constraint: We tuned the KL coefficient to maintain a small, constant KL divergence during training, keeping the model close to its pretrained weight. This allows for task-specific reward optimization without harming general performance. As shown in Table 2, removing the KL constraint boosts performance on Compositional Image Generation and Visual Text Rendering but causes a sharp decline in image quality and preference scores on DrawBench. In contrast, a properly tuned KL constraint preserves quality while maintaining similar gains on task-specific metrics. In the Human Preference Alignment task, dropping the KL constraint doesn’t hurt image quality—likely due to PickScore overlap with evaluation metrics—but causes a collapse in visual diversity. Outputs converge to a single style, with different seeds producing nearly identical results. KL regularization prevents this collapse and preserves diversity. See Figure 7 in the Appendix C.2 for learning curves with and without KL, and Figure 4 for examples.
Figure 4: Effect of KL Regularization. The KL penalty effectively suppresses reward hacking, preventing potential Quality Degradation (for GenEval and OCR) and Diversity Decline (for PickScore).
Effect of Denoising Reduction
Figure 5 (a) highlights Denoising Reduction’s significant impact on accelerating training. To explore how different timesteps affect optimization, these experiments were conducted without the KL constraint. Reducing data collection timesteps from $40$ to $10$ achieved over a 4×4$\times$4 × speedup across all three tasks, without impacting final reward. While further tests showed this reduction didn’t consistently improve speed (and occasionally slowed some tasks), we standardized on $10$ timesteps for subsequent experiments, prioritizing the maintained reward. For the other two tasks, learning curves of reward versus training time are presented in Figure 6 in the Appendix C.1.
Effect of Noise Level
Higher $\sigma_{t}$ in the SDE boosts image diversity and exploration, vital for RL training. We control this exploration with a noise level $a$ (Eq. 12). Ablation studies (Figure 5 (b)) show $a$’s impact on performance. Small $a$ (e.g., 0.1) limits exploration and slows reward improvement. Increasing $a$ (up to 0.7) enhances exploration and accelerates reward gains. Beyond this, further increases (e.g., $a=0.7\rightarrow 1.0$) yield no extra benefit, as exploration is sufficient. However, too much noise degrades image quality, leading to zero reward and failed training. We advise using the highest noise level that maintains good image quality.
(a). Effect of Denoising Reduction on GenEval
(b). Effect of Noise Level Ablation on the OCR
Figure 5: Ablation studies on our critical design choices. (a) Denoising Reduction: Fewer denoising steps accelerate convergence and yield similar performance. (b) Noise Level: Moderate noise level ($a=0.7$) maximises OCR accuracy, while too little noise hampers exploration.
Generalization Analysis
Flow-GRPO demonstrates strong generalization on unseen scenarios from GenEval (Table 4). Specifically, it captures object number, color, and spatial relations, generalizing well to unseen object classes. It also effectively controls object count, generalizing from training on $2-4$ objects to generate $5-6$ objects. Furthermore, Table 3 shows Flow-GRPO achieves significant gains on T2I-CompBench++. This comprehensive benchmark for open-world compositional T2I generation features object classes and relationships substantially different from our model’s GenEval-style training data.
Table 3: T2I-CompBench++ Result. This evaluation uses the same model presented in Table 1, which was trained on the GenEval-generated dataset. The best score is in blue.
Model Color Shape Texture 2D-Spatial 3D-Spatial Numeracy Non-Spatial
Janus-Pro-7B [63] 0.5145 0.3323 0.4069 0.1566 0.2753 0.4406 0.3137
EMU3 [61] 0.7913 0.5846 0.7422
FLUX.1 Dev [5] 0.7407 0.5718 0.6922 0.2863 0.3866 0.6185 0.3127
SD3.5-M 0.7994 0.5669 0.7338 0.2850 0.3739 0.5927 0.3146
SD3.5-M+Flow-GRPO 0.8379 0.6130 0.7236 0.5447 0.4471 0.6752 0.3195
Table 4: Flow-GRPO demonstrates strong generalization. Unseen Objects: Trained on 60 object classes, evaluated on 20 unseen classes. Unseen Counting: Trained to render 2, 3, or 4 objects, was evaluated on rendering 5 or 6.
Method Unseen Objects Unseen Counting
Overall Single Obj. Two Obj. Counting Colors Position Attr. Binding Counting
SD3.5-M 0.64 0.96 0.73 0.53 0.87 0.26 0.47 0.13
SD3.5-M+Flow-GRPO 0.90 1.00 0.94 0.86 0.97 0.84 0.77 0.48
Limitations & Future Work.
While this work focuses on T2I tasks, Flow-GRPO’s potential applicability to video generation [25, 27] raises several questions for future research: (1) Reward Design for Video: Defining effective reward models for video generation is crucial. Simple heuristics like object detectors or trackers can promote physical realism and temporal consistency, but more sophisticated models are needed. (2) Balancing Multiple Rewards: Effective video generation requires optimizing multiple objectives (e.g., realism, smoothness, coherence). Balancing these often conflicting signals is challenging and demands careful tuning. (3) Scalability: Video generation is significantly more resource-intensive than T2I. Scaling Flow-GRPO for video thus demands more efficient data collection and training pipelines.
Appendix A Mathematical Derivations for Stochastic Sampling using Flow Models
We present a detailed proof here. To compute $p_{\theta}({\bm{x}}_{t-1}\mid{\bm{x}}_{t},{\bm{c}})$ in Equation 6 during forward sampling, we adapt flow models to a stochastic differential equation (SDE). While flow models normally follow a deterministic ODE:
$\mathrm{d}{\bm{x}}_{t}={\bm{v}}_{t}\mathrm{d}t$ (14)
We consider its stochastic counterpart. Inspired by the derivation from SDE to its probability flow ODE in SGMs [23], we aim to construct a SDE with specific drift and diffusion coefficients so that its marginal distribution matches that of Eq. 14. We begin with the generic form of SDE:
$\mathrm{d}{\bm{x}}_{t}=f_{\text{SDE}}({\bm{x}}_{t},t)\mathrm{d}t+\sigma_{t}\mathrm{d}{\bm{w}},$ (15)
Its marginal probability density $p_{t}({\bm{x}})$ evolves according to the Fokker–Planck equation [69], i.e.,
$\partial_{t}p_{t}(x)=-\nabla\cdot[f_{\text{SDE}}({\bm{x}}_{t},t)p_{t}({\bm{x}})]+\frac{1}{2}\nabla^{2}[\sigma_{t}^{2}p_{t}({\bm{x}})]$ (16)
Similarly, the marginal probability density associated with Eq. 14 evolves:
$\partial_{t}p_{t}({\bm{x}})$ $=-\nabla\cdot[{\bm{v}}_{t}({\bm{x}}_{t},t)p_{t}({\bm{x}})]$ (17)
To ensure that the stochastic process shares the same marginal distribution as the ODE, we impose:
$-\nabla\cdot[f_{\text{SDE}}p_{t}({\bm{x}})]+\frac{1}{2}\nabla^{2}[\sigma_{t}^{2}p_{t}({\bm{x}})]=-\nabla\cdot[{\bm{v}}_{t}p_{t}({\bm{x}})]$ (18)
Observing that
$\nabla^{2}[\sigma_{t}^{2}p_{t}({\bm{x}})]$ $=\sigma_{t}^{2}\nabla^{2}p_{t}({\bm{x}})$ (19)
$=\sigma_{t}^{2}\nabla\cdot(\nabla p_{t}({\bm{x}}))$
$=\sigma_{t}^{2}\nabla\cdot(p_{t}({\bm{x}})\nabla\log p_{t}({\bm{x}}))$
Substituting Eq. 19 to Eq. 18, we arrive at the drift coefficients of the target SDE:
$f_{\text{SDE}}={\bm{v}}_{t}-\frac{1}{2}\sigma_{t}^{2}\nabla\log p_{t}({\bm{x}})$ (20)
Hence we can rewrite the reverse-time SDE in Eq. 15 as:
$\mathrm{d}{\bm{x}}_{t}=\left({\bm{v}}_{t}({\bm{x}}_{t})-\frac{\sigma_{t}^{2}}{2}\nabla\log p_{t}({\bm{x}}_{t})\right)\mathrm{d}t+\sigma_{t}\mathrm{d}{\bm{w}},$ (21)
where $d{\bm{w}}$ denotes Wiener process increments and $\sigma_{t}$ is the diffusion coefficient that controls the level of stachasticity during sampling.
Once the score $\nabla\log p_{t}({\bm{x}}_{t})$ is available, the stochastic process can be sampled directly. While the Flow-Matching framework, this score is tied to the velocity field ${\bm{v}}_{t}$.
Specifically, let $\dot{\alpha_{t}}\equiv\partial\alpha_{t}/\partial t$. All expectations are over ${\bm{x}}_{0}\sim X_{0}$ and ${\bm{x}}_{1}\sim\mathcal{N}(0,{\bm{I}})$, where $X_{0}$ is the data distribution.
For the linear interpolation ${\bm{x}}_{t}=\alpha_{t}{\bm{x}}_{0}+\beta_{t}{\bm{x}}_{1}$, we have:
$p_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})=\mathcal{N}\left({\bm{x}}_{t}\mid\alpha_{t}{\bm{x}}_{0},\beta_{t}^{2}{\bm{I}}\right),$ (22)
yielding the conditional score:
$\nabla\log p_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})=-\frac{{\bm{x}}_{t}-\alpha_{t}{\bm{x}}_{0}}{\beta_{t}^{2}}=-\frac{{\bm{x}}_{1}}{\beta_{t}}.$ (23)
The marginal score becomes:
$\nabla\log p_{t}({\bm{x}}_{t})$ $=\mathbb{E}\left[\nabla\log p_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})\mid{\bm{x}}_{t}\right]$
$=-\frac{1}{\beta_{t}}\mathbb{E}[{\bm{x}}_{1}\mid{\bm{x}}_{t}].$ (24)
For the velocity field ${\bm{v}}_{t}({\bm{x}}_{t})$, we derive:
$\begin{split}{\bm{v}}_{t}({\bm{x}})&=\mathbb{E}\left[\dot{\alpha_{t}}{\bm{x}}_{0}+\dot{\beta_{t}}{\bm{x}}_{1}\mid{\bm{x}}_{t}={\bm{x}}\right]\\ &=\dot{\alpha_{t}}\mathbb{E}[{\bm{x}}_{0}\mid{\bm{x}}_{t}={\bm{x}}]+\dot{\beta_{t}}\mathbb{E}[{\bm{x}}_{1}\mid{\bm{x}}_{t}={\bm{x}}]\\ &=\dot{\alpha_{t}}\mathbb{E}\left[\frac{{\bm{x}}_{t}-\beta_{t}{\bm{x}}_{1}}{\alpha_{t}}\mid{\bm{x}}_{t}={\bm{x}}\right]+\dot{\beta_{t}}\mathbb{E}[{\bm{x}}_{1}\mid{\bm{x}}_{t}={\bm{x}}]\\ &=\frac{\dot{\alpha_{t}}}{\alpha_{t}}{\bm{x}}-\frac{\dot{\alpha_{t}}\beta_{t}}{\alpha_{t}}\mathbb{E}[{\bm{x}}_{1}\mid{\bm{x}}_{t}={\bm{x}}]+\dot{\beta_{t}}\mathbb{E}[{\bm{x}}_{1}\mid{\bm{x}}_{t}={\bm{x}}]\\ &=\frac{\dot{\alpha_{t}}}{\alpha_{t}}{\bm{x}}-\left(\dot{\beta_{t}}\beta_{t}-\frac{\dot{\alpha_{t}}\beta_{t}^{2}}{\alpha_{t}}\right)\nabla\log p_{t}({\bm{x}}),\end{split}$ (25)
Substituting $\alpha_{t}=1-t$ and $\beta_{t}=t$ simplifies Equation 25 to:
${\bm{v}}_{t}({\bm{x}})=-\frac{{\bm{x}}}{1-t}-\frac{t}{1-t}\nabla\log p_{t}({\bm{x}}).$ (26)
Solving for the score yields:
$\nabla\log p_{t}({\bm{x}})=-\frac{{\bm{x}}}{t}-\frac{1-t}{t}{\bm{v}}_{t}({\bm{x}}).$ (27)
Substituting Equation 27 into 21 gives the final SDE:
$\mathrm{d}{\bm{x}}_{t}=\left[{\bm{v}}_{t}({\bm{x}}_{t})+\frac{\sigma_{t}^{2}}{2t}\left({\bm{x}}_{t}+(1-t){\bm{v}}_{t}({\bm{x}}_{t})\right)\right]\mathrm{d}t+\sigma_{t}\mathrm{d}{\bm{w}}.$ (28)
Applying Euler-Maruyama discretization yields the update rule:
${\bm{x}}_{t+\Delta t}={\bm{x}}_{t}+\left[{\bm{v}}_{\theta}({\bm{x}}_{t},t)+\frac{\sigma_{t}^{2}}{2t}\big{(}{\bm{x}}_{t}+(1-t){\bm{v}}_{\theta}({\bm{x}}_{t},t)\big{)}\right]\Delta t+\sigma_{t}\sqrt{\Delta t}\,\epsilon,$ (29)
where $\epsilon\sim\mathcal{N}(0,{\bm{I}})$ injects stochasticity.
Appendix B Further Details on the Experimental Setup
B.1 Model Specification
The following table lists the base model and the reward models and their corresponding links.
B.2 Hyperparameters Specification
Except for $\beta$, GRPO hyperparameters are fixed across tasks. We use a sampling timestep $T=10$ and an evaluation timestep $T=40$. Other settings include a group size $G=24$, an noise level $a=0.7$ and an image resolution of 512. The KL ratio $\beta$ is set to 0.004 for GenEval and Text Rendering, and 0.001 for Pickscore. We use Lora with $\alpha=64$ and $r=32$.
B.2.1 Compute Resources Specification
We train our model using 24 NVIDIA A800 GPUs. The learning curves in Appendix C.2 provide details on the specific GPU hours.
Appendix C Extended Experimental Results
C.1 Effect of Denoising Reduction
We show the extended Denoising Reduction ablations of Visual Text Rendering and Human Preference Alignment tasks in Figure 6.
(a). Visual Text Rendering
(b). Human Preference Alignment
Figure 6: Effect of Denoising Reduction
C.2 Learning Curves with or without KL
Figure 7 shows learning curves for three tasks, with and without KL. These results emphasize that KL regularization is not empirically equivalent to early stopping. Adding appropriate KL can achieve the same high reward as the KL-free version and maintain image quality, though it requires longer training.
(a). Compositional Image Generation
(b). Visual Text Rendering
(c). Human Preference Alignment
Figure 7: Learning Curves with and without KL. KL penalty slows early training yet effectively suppresses reward hacking.
C.3 Qualitative Results
In addition to the GenEval reward comparisons in Figure 3, Figures 8 and 9 show qualitative comparisons between SD-3.5-M and SD-3.5-M + Flow-GRPO trained with PickScore and OCR rewards, respectively.
Figure 8: Qualitative comparison between the SD-3.5-M and SD-3.5-M + Flow-GRPO trained with PickScore reward.
Figure 9: Qualitative comparison between the SD-3.5-M and SD-3.5-M + Flow-GRPO trained with OCR reward.